11. Which term of the A.P. 3, 15, 27, 39,.. will be 132 more than its 54th term?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given A.P. is 3, 15, 27, 39, … first term, a = 3 common difference, d = a2 − a1 = 15 − 3 = 12 We know that, an = a+(n−1)d Therefore, a54 = a+(54−1)d ⇒3+(53)(12) ⇒3+636 = 639 a54 = 639 We have to find the term of this A.P. which is 132 more than a54, i.e.771. Let nth term be 771. an = a+(n−1)d 771 = 3+(n −1)12 768 = (n−1)12 (n −1) = 64 n = 65 Therefore, 65th term was 132 more than 54th term. Or another method is; Let nth term be 132 more than 54th term. n = 54 + 132/2 = 54 + 11 = 65th term